OpenStudy (anonymous):

please help Samantha knows that she needs $22,000 for a 10% down payment on a house she can afford. She found an investment that earns 2.75% interest compounding monthly. How much should she put in the account now, rounded to the nearest dollar, to ensure she has the down payment amount in 5 years? $4,320 $19,177 $13,371 $5,649

OpenStudy (anonymous):

@ivettef365 @mathmate @mathstudent55

OpenStudy (anonymous):

i need help solving it @mathstudent55

OpenStudy (anonymous):

Ok now what?

OpenStudy (mathstudent55):

Sorry. I misread the problem. It's easier than I thought.

OpenStudy (anonymous):

Lol ok omg I was getting so confused.

OpenStudy (mathstudent55):

\(F = P(1 + i)^n \) where F = future value ($22,000) P = present value (unknown) n = number of compounding periods (60) i = interest per period expressed as decimal (0.0275/12)

OpenStudy (anonymous):

Ok can you help me solve??

OpenStudy (mathstudent55):

\($22,000 = P(1 + \frac{0.0275}{12})^{60}\)

OpenStudy (mathstudent55):

\($22,000 = P(1.14722) \) \(P = \dfrac{$22,000}{1.14722} \) \(P = $19,177 \)